A light pointer fixed to one prong of a tuning fork touches a vertical plate. the fork is set vibrating at a frequency of 56 Hz and allowed to free fall . how many complete oscillatons are counted when platye falls at10 cm.

Grade:9

1 Answers

Arun
6 years ago
Dear Mukesh
 
t = \LARGE \sqrt(2*0.1)/9.8
t = (1/7) sec
 
Frequency is 56 Hz. Which means in 1 sec it makes 56 vibrations.
Hence in 1/7 sec it will make = 56/7 = 8 counts
 
 
Regards
Arun(askIITians forum expert)